# Exergetic Port-Hamiltonian Systems Modeling Language

### v1

@@ Abstract
Mathematical modeling of real-world physical systems requires the consistent combination of a multitude of physical laws and phenomenological models. This challenging task can be greatly simplified by hierarchically decomposing systems. Moreover, the use of diagrams for expressing such decompositions helps make the process more intuitive and facilitates communication, even with non-experts. As an important requirement, models have to respect fundamental physical laws such as the first and the second law of thermodynamics. While some existing modeling frameworks can make such guarantees based on structural properties of their models, they lack a formal graphical syntax. We present a compositional and thermodynamically consistent modeling language with a graphical syntax. As its semantics, port-Hamiltonian systems are endowed with further structural properties and a fixed physical interpretation such that thermodynamic consistency is ensured in a way that is closely related to the GENERIC framework. While port-Hamiltonian systems are inspired by graphical modeling with bond graphs, neither the link between the two, nor bond graphs themselves, can be easily formalized. In contrast, our syntax is based on a refinement of the well-studied operad of undirected wiring diagrams. The language effectively decouples the construction of complex models via the graphical syntax from physical concerns, which are dealt with only at the level of primitive subsystems that represent elementary physical behaviors. As a consequence, reuse of models and substitution of their parts becomes possible. Finally, by construction, systems interact by exchanging exergy, i.e. energy that is available for doing work, so the language is particularly well suited for thermodynamic analysis and optimization.
@@
applied category theory, bond graphs, classical electromagnetism, classical mechanics, compositionality, exergy, GENERIC, graphical syntax, local thermodynamic equilibrium, multiphysics, nonequilibrium thermodynamics, port-Hamiltonian systems
@@ 1 Introduction
In this introduction we want to answer the following questions:
@@
In addition, we briefly summarize key ideas of nonequilibrium thermodynamics and we discuss related work that inspired our developments. The introduction finishes with an outline of the main part of the paper.
@@ 1.1 What?
EPHS provide a compositional and thermodynamically consistent language for expressing mathematical models of multiphysical systems at macroscopic scales. Such systems may combine
@@
The EPHS language is termed compositional since it enables a completely modular and hierarchical specification of models based on its simple graphical syntax, which is discussed in more detail in the next subsection.
If one flattens any hierarchical nesting of systems, one finds that every system is determined by a power-preserving interconnection of primitive subsystems. There are three main classes of primitive systems:
@@
    Systems that represent energy storage are characterized by a function which yields the stored energy for each possible state of the primitive system.
@@
    Systems that represent reversible energy exchange are characterized by a so-called Dirac structure, as known e.g. from (constrained) mechanical systems Courant (1990).
@@
Since both classical mechanics and electromagnetism can be expressed using Dirac structures, item 2. corresponds to (a) and (b) above, while item 3. corresponds to (c). Structural properties of the primitive systems guarantee that all models respect the first and the second law of thermodynamics, Onsager’s reciprocal relations, as well as further conservation laws (e.g. for mass). Section 1.4 aims for a very brief introduction to nonequilibrium thermodynamics.
@@ 1.2 How?
Figure 1: Graphical representation of an interconnection pattern with different syntactic elements shown in a different color. The outer box shown in orange represents the interface of the resulting composite system. The two protruding lines correspond to its ports called $\mathtt{f}$ and $\mathtt{b_{k}}$. The two inner boxes shown in red represent (the interface of) the subsystems called $\mathtt{ke}$ (storage of kinetic energy) and $\mathtt{sa}$ (self-advection of kinetic energy). Both have ports called $\mathtt{p}$ (momentum) and $\mathtt{m}$ (mass), while $\mathtt{sa}$ has another port called $\mathtt{b_{k}}$ (advection of kinetic energy across the boundary of the spatial domain). The port combinator shown in purple combines the ports $\mathtt{p}$ (fluid momentum) and $\mathtt{m}$ (fluid mass) into the multiport $\mathtt{f}$ (fluid). The black dots represent junctions where power is exchanged among the connected ports. The dashed line indicates that port $\mathtt{m}$ of the port combinator does not partake in power exchange, but merely shares information about the state (fluid mass).
Expressions in the graphical syntax are called interconnection patterns because a pattern specifies how a composite system can be formed through a power-preserving interconnection of subsystems. Figure 1 shows an example from fluid dynamics.
@@
An interconnection pattern should be thought of as a function which, for every round inner box, takes as input a system with matching interface and returns a system whose interface matches the outer box. For instance, the pattern shown in Figure 1 represents a function which takes two subsystems whose interfaces are determined by the ports of the boxes $\mathtt{ke}$ and $\mathtt{sa}$, and it returns a composite system whose interface is determined by the ports of the outer box. In the graphical representation, we often draw inner boxes filled with a color that indicates the nature of the subsystems that we consider as inputs. In Figure 1, $\mathtt{ke}$ has a blue filling as it represents energy storage, i.e. the fluid’s kinetic energy, while $\mathtt{sa}$ has a green filling as it represents reversible energy exchange, i.e. the advection of momentum and mass by the fluid flow. Written on top of the outer box, kin is an identifier akin to a function name and the text in parenthesis is just a short description. The port combinator and the corresponding multiport $\mathtt{f}$ are syntactic sugar that allows the composite system to be connected to other systems without drawing parallel connections.
@@
Figure 2: Hierarchical specification of an electro-magneto-hydrodynamics model: The fluid model f shown in (b) comprises the kinetic energy system kin shown in (a) and four further models which are not shown, namely an internal energy system int, a thermal conduction model th, and volume / shear viscosity models vol / shr. The electro-magneto-hydrodynamics model emhd shown in (c) comprises the fluid model f and three further models which are not shown, namely an electro-magnetic system em, an electric conduction model el, and a model for the electro-kinetic coupling ek.
Figure 2 demonstrates the hierarchical specification of a model for electro-magneto-hydrodynamics. For simplicity, a spatial domain with empty or isolated boundary is assumed such that only the in-domain interconnection of the different models needs to be dealt with. The yellow filling of the inner boxes in Figures 1(b) and 1(c) indicates that another composite system, which in turn is specified with its own interconnection pattern, is supplied as input. Recursively substituting into all yellow inner boxes the respective interconnection patterns would yield an equivalent model specification without hierarchical decomposition. The composition of interconnection patterns through substitution is uniquely defined as long as interfaces match. Similarly, one can add a level of nesting by refactoring a number of subsystems into a reusable composite subsystem. In the fluid model, one could for instance factor out the kinetic energy system filling the box $\mathtt{kin}$ and the internal energy subsystem filling the box $\mathtt{int}$ into a reusable ideal compressible fluid model.
@@ 1.3 Why?
EPHS provide a formal and intuitive modeling language suited for humans and software tools alike. As shown above, increasingly complex models can be assembled from simpler and ultimately primitive parts. This fosters the reusability of models and encourages the conceptual separation of different levels of detail. Beyond the modeling process, this provides value in interdisciplinary communication and education. Some modeling errors can be completely avoided due to typed interfaces (given by ports with fixed physical quantities) and the built-in thermodynamic consistency. Since interconnection patterns can be used as exergy flow diagrams, EPHS models are particularly useful for thermodynamic analysis and optimization. The compositional nature of EPHS could enable a scalable approach to model transformations, such as discretizations, and also to the integration of scientific machine learning techniques in first-principles-based physical modeling.
1.4 Nonequilibrium thermodynamics in a tiny nutshell
@@
Most physical systems are characterized by a huge number of degrees of freedom but keeping track of them is usually neither possible nor useful. Nonequilibrium thermodynamics aims to provide a framework for coarse-grained modeling, meaning that fast dynamics at small scales are not resolved and a quantity called entropy is introduced to account for the missing information, see e.g. Pavelka et al. (2018). Coarse-graining is possible because, despite of fluctuations at the microscopic scale, one can eventually observe non-fluctuating properties at a more macroscopic scale; at least this is the case for isolated systems. For instance, temperature is a property that does not appear at the microscopic scale. Thermodynamic properties are classified as extensive when they are proportional to the size of the system (e.g. entropy, volume, mass) or as intensive when they do not depend on the size of the system (e.g. temperature, pressure, chemical potential), see e.g. Callen (1985)
@@
A system is said to be in local thermodynamic equilibrium if for every material point there exists a neighborhood large enough to observe thermodynamic properties, yet small enough for the properties to be non-varying inside the neighborhood, see e.g. Öttinger (2005). The size of these neighborhoods presents a macroscopic scale at which the dynamics of the thermodynamic properties can be observed and also modeled.
@@
The first law of thermodynamics just restates a well-known fact from mechanics, namely that energy is a conserved quantity. However, instead of considering directly the mechanical and electromagnetic energy associated to microscopic degrees of freedom, a thermodynamic model with local equilibrium resorts to a phenomenological description called internal energy, which is considered to depend on a number of extensive properties, including entropy, and possibly further mechanical or electromagnetic state variables which are observable on the considered macroscopic scale. These quantities then serve as the state variables of a coarse-grained description. Moreover, temperature is the derivative of the internal energy with respect to entropy. Similarly, pressure is given by the derivative with respect to volume and the chemical potential is the derivative with respect to mass, see e.g. Callen (1985).
@@
Classical mechanics and electromagnetism exhibit reversible dynamics, meaning that one could in principle arrive again at the initial condition by reversing the direction of time (although for many systems this would be impossible due to chaos). Incomplete knowledge about the microscopic state immediately leads to irreversible dynamics and an irreversible dynamics in turn leads to a growth of entropy, as information (e.g. about the initial state of the system) is lost in the process. This is reflected by the second law of thermodynamics which states that the entropy of an isolated system never decreases and that the system evolves towards a state of maximum entropy, see e.g. Öttinger (2005). In other words, the uncertainty about the microscopic state can only grow, not decrease.
@@
For models with local equilibrium, the irreversible dynamics can be expressed in terms of so-called thermodynamic fluxes that determine instantaneous changes of coarse-grained state variables. The fluxes are expressed as a function of thermodynamic forces, which are given by differences in intensive quantities, see e.g. Öttinger (2005); de Groot and Mazur (1984). For instance, temperature differences cause heat flux, according to Fourier’s law, and differences in chemical potential cause a diffusive mass flux, according to Fick’s law. Further, one may observe cross effects such as a diffusive mass flux caused by temperature differences (Soret effect) and likewise a heat flux caused by differences in chemical potential (Dufour effect). Onsager’s reciprocal relations state that the relation between fluxes and forces must possess a certain symmetry Onsager (1931). For instance, the ratio of the mass flux caused by the Soret effect and the respective temperature difference is equal to the ratio of the heat flux caused by the Dufour effect and the respective difference in chemical potential.
@@ 1.5 Related work
Figure 3: Informal Venn diagram illustrating that EPHS integrates ideas from graphical and energy-based modeling of physical systems with bond graphs, the GENERIC framework for nonequilibrium thermodynamics, the port-Hamiltonian theory of open dissipative systems (PHS), and applied category theory (ACT) research on the formalization of graphical languages and compositional dynamical systems.
@@ 1.5.1 Port-Hamiltonian systems
Apparent from their name, EPHS have a lot in common with Port-Hamiltonian Systems (PHS). PHS were originally inspired by relating the dynamical systems defined by generalized bond graphs Breedveld (1982) (i.e. bond graphs using a single kind of energy storage elements together with gyrator elements) to Hamiltonian mechanics Maschke (1991). While many authors make use of explicit input-output port-Hamiltonian systems, PHS constitute an acausal systems theory, inspired also by the behavioral approach of Willems Willems (1974, 2007). The dynamics of networks comprising energy storing elements and reversible coupling elements between them (such as generalized transformers and gyrators) can be formalized, from a geometric viewpoint, as a Hamiltonian system whose Poisson structure reflects the network topology Maschke et al. (1992). The generalization to networks with algebraic constraints and externally controlled ports is achieved by using Dirac structures, which unify Poisson structures and presymplectic structures van der Schaft and Maschke (1995). By specifying a resistive relation on external port variables, the implicit Hamiltonian dynamics can be combined with a dissipative gradient dynamics, allowing for network elements that represent e.g. dampers and resistors. The key properties of PHS are firstly a power-balance equation which states that the sum of the stored power and the dissipated power is equal to the supplied power (passivity) and secondly the possibility to combine port-Hamiltonian systems by composition of their Dirac structures with an interconnecting Dirac structure, see for instance van der Schaft and Jeltsema (2014).
While the dynamical systems defined by EPHS models are essentially PHS, they have further structural properties and a fixed physical interpretation. PHS in general have no uniform thermodynamic interpretation and they lack a formal graphical syntax to express their interconnection. In contrast to EPHS, PHS also lack a built-in mechanism for referring to subsystems and ports.
1.5.2 GENERIC framework
The general equation for reversible-irreversible coupling (GENERIC) provides a framework for combining reversible and irreversible dynamics in a way that respects fundamental thermodynamic principles, namely the first law, the second law and the reciprocal relations discovered by Onsager Grmela and Öttinger (1997). The reversible part is defined by an entropy-conserving Hamiltonian dynamics, while the irreversible part is given by an energy-conserving gradient flow, see e.g. Öttinger (2005). We note that the same holds for closely-related earlier developments in this direction, including metriplectic systems, see Kaufman (1984); Grmela (1984a, b); Morrison (1984). While a large class of well-known dynamical systems from physics can be identified as instances of the GENERIC, see e.g. Pavelka et al. (2018), the framework is especially helpful for the development of new thermodynamic models.
The combination of Hamiltonian and gradient dynamics in both PHS and GENERIC systems led to the initial observation of a relationship between the two frameworks in Badlyan et al. (2018). GENERIC systems with local thermodynamic equilibrium can be seen as port-Hamiltonian systems, where the storage function corresponds to the system’s exergy. This function yields the maximum amount of work that can be extracted by an ideal (i.e. purely reversible) process which may arbitrarily interact with the system’s storage components and a given reference environment Lohmayer et al. (2021).
The dynamical systems defined by (isolated) EPHS models are basically GENERIC systems, since their built-in thermodynamic consistency and associated physical interpretation follow from the same ideas. Beyond the aspect of compositionality, there are two main differences: 1. GENERIC systems are more general in the sense of including systems beyond local thermodynamics equilibrium (where the concept of temperature does not apply). 2. The reversible part of EPHS is defined using Dirac structures, rather than less general Poisson structures (which cannot express kinematic constraints). This is crucial e.g. for modeling multibody systems and electrical circuits.
1.5.3 Bond graph modeling
Bond graph modeling is based on a graphical notation for representing dynamical systems and in particular multi-domain models involving mechanical and electrical elements Paynter (1961); Borutzky (2010). The main idea is that elements interact via bonds that represent links for energy exchange. Associated to each bond, there are two dynamic variables called flow and effort, whose product (or duality pairing) gives the exchanged power. Bonds can be connected via two types of junctions, namely 00-junctions, where the flow variables of all connected ports balance, with all effort variables being equal, and dually defined $1$-junctions, where the effort variables balance, with all flow variables being equal. In both cases, the net power at a junction is zero.
While PHS can be seen as a geometric formalization of the dynamical systems defined by generalized bond graphs, the algebraic formalization of the bond graph notation itself is not straightforward, when compared to e.g. standard multigraphs, see Birkett and Roe (1989); Coya (2017). The inherent complexity of the notation is evident also in practice, since the direction in which exchange of power is counted positive and sometimes also the question which variable is flow and which is effort have to be decided arbitrarily, bond by bond. While bond graphs are frequently used in the literature to illustrate PHS, their use as a formal syntax for model specification is not established. Previous work has focused on algorithms for converting certain classes of bond graphs to port-Hamiltonian systems, see e.g. Pfeifer et al. (2020).
In contrast to bond graph notation, the EPHS syntax only has 00-junctions and every interconnection pattern has an outer box. As demonstrated by various applications, a single type of junctions is sufficient and moreover the simplification entails crucial advantages: Interconnection patterns are easier to interpret, in particular due to the direct correspondence between junctions and energy domains. This correspondence arises from the inherent distinction between the roles of flows and efforts, a feature that is lacking in bond graphs. Further, the direction in which exchange of power is counted positive is predetermined in accordance with the convention used in thermodynamics, i.e. energy supplied to a system has a positive sign. Last but not least, interconnection patterns can be hierarchically composed and decomposed in a natural way.

### v2

@@ Abstract
Mathematical modeling of real-world physical systems requires the consistent combination of a multitude of physical laws and phenomenological models. This challenging task can be greatly simplified by hierarchically decomposing systems into ultimately simple components. Moreover, the use of diagrams for expressing the decomposition helps make the process more intuitive and facilitates communication, even with non-experts. As an important requirement, models have to respect fundamental physical laws such as the first and the second law of thermodynamics. While some existing modeling frameworks make such guarantees based on structural properties of their models, they lack a formal graphical syntax. We present a compositional and thermodynamically consistent modeling language with a graphical syntax. In terms of its semantics, we essentially endow port-Hamiltonian systems with additional structural properties and a fixed physical interpretation, ensuring thermodynamic consistency in a manner closely related to the metriplectic or GENERIC formalism. While port-Hamiltonian systems are inspired by graphical modeling with bond graphs, neither the link between the two, nor bond graphs themselves, can be easily formalized. In contrast, our syntax is based on a refinement of the well-studied operad of undirected wiring diagrams. By combining a compositional, graphical syntax with an energy-based, thermodynamic approach, the presented modeling language simplifies the understanding, reuse, and modification of complex physical models.
@@
bond graphs , compositionality , GENERIC , metriplectic structure , nonequilibrium thermodynamics , undirected wiring diagrams
@@ 1 Introduction
In this introduction we mainly want to answer the following questions:
@@
In addition, we briefly summarize key ideas of nonequilibrium thermodynamics, and we discuss related work that inspired our developments. The introduction finishes with an outline of the main part of the paper.
@@ 1.1 What?
Exergetic Port-Hamiltonian Systems (EPHS) provide a compositional and thermodynamically consistent language for expressing mathematical models of multiphysical systems at macroscopic scales. Such systems may combine
@@
The EPHS language is termed compositional because it enables a fully modular and hierarchical approach to model specification through its simple, graphical syntax, as elaborated in the next subsection.
Flattening any hierarchical nesting reveals that every EPHS model is fundamentally defined by a power-preserving interconnection of primitive subsystems. There are three main classes of primitive EPHS:
@@
    Systems that represent energy storage are characterized by a function which yields the stored energy for each state of the primitive system.
@@
    Systems that represent reversible energy exchange are characterized by a so-called Dirac structure [1], as known e.g. from (constrained) mechanical systems.
@@
Since both classical mechanics and electromagnetism can be expressed using Dirac structures, item 2. corresponds to (a) and (b) above, while item 3. corresponds to (c). Structural properties of the primitive systems and their interconnection guarantee that all models respect the first and the second law of thermodynamics, as well as Onsager’s reciprocal relations.
Inspired by the contribution in [2], previous work [3] has explored EPHS as a systematic approach to thermodynamic modeling based on port-Hamiltonian systems (see, e.g., [4]). By interpreting the ‘Hamiltonian’ as an exergy storage function, a stricter version of port-Hamiltonian structure can be formulated, ensuring a thermodynamically consistent reversible-irreversible splitting of the dynamics, as known from the metriplectic or GENERIC formalism (see, e.g., [5, 6]). The exergy storage function represents the theoretically ‘useful’ amount of energy, i.e., the maximum amount of work that can be extracted from a thermodynamic system using an ideal device, such as a Carnot engine. This ideal device has access not only to all forms of energy stored in the system but also to a (reference) environment, primarily characterized by a heat bath with constant temperature.
The present work formalizes EPHS as a compositional modeling language with a simple, graphical syntax inspired by bond graphs (see, e.g., [7, 8]). This approach enables the construction of complex models by composing them from simpler subsystems, which can themselves be constructed from more basic subsystems, and so on. We draw on and extend ideas form (applied) category theory (see, e.g., [9, 10]) to propose a simple mathematical framework suitable for computer implementation. In principle, the graphical syntax and the associated compositional framework could also be adopted without the parts that ensure thermodynamic consistency. While this would relax certain conditions imposed on the primitive systems, it would also weaken the correspondence between the remaining structure and its physical interpretation.
@@ 1.2 How?
Figure 1: Graphical display of an interconnection pattern with different syntactic elements shown in different colors. The outer box shown in purple represents the interface of the resulting composite system. The three protruding lines represent its ports called $\mathtt{f.p}$, $\mathtt{f.m}$ and $\mathtt{b_{k}}$. The two inner boxes shown in red represent (the interface of) the subsystems called $\mathtt{ke}$ (storage of kinetic energy) and $\mathtt{sa}$ (self-advection of kinetic energy). Both have ports called $\mathtt{p}$ (momentum) and $\mathtt{m}$ (mass). Box $\mathtt{sa}$ has another port called $\mathtt{b_{k}}$ (advection of kinetic energy across the boundary of the spatial domain). The common prefix $\mathtt{f}$ (fluid) of the outer ports $\mathtt{f.p}$ (fluid momentum) and $\mathtt{f.m}$ (fluid mass) combines them into a multiport. The black dots represent junctions where power is exchanged among the connected ports. The dashed line indicates that only information about the state (fluid mass) is exchanged via the outer port $\mathtt{f.m}$, with no energy transfer occurring.
Expressions in the graphical syntax are called interconnection patterns, since such a pattern specifies how a composite system is formed through a power-preserving interconnection of subsystems. Figure 1 shows an example from fluid dynamics, see [11].
@@
We think of an interconnection pattern as a function. For each inner box, it takes as input a system with matching interface and it outputs a system whose interface matches the outer box. For instance, the pattern shown in Figure 1 represents a function that takes two subsystems, whose interfaces are determined by the ports of the boxes $\mathtt{ke}$ and $\mathtt{sa}$, and it combines them into a composite system, whose interface is given by the ports of the outer box. We draw inner boxes filled with a color to indicate the nature of the subsystems that we consider as inputs. In Figure 1, box $\mathtt{ke}$ has a blue filling, as it represents energy storage (the fluid’s kinetic energy). Box $\mathtt{sa}$ has a green filling, as it represents reversible energy exchange (advection of momentum and mass). Written on top of the outer box, kin is an identifier akin to a function name and the text in parentheses is just a short description. The multiport $\mathtt{f}$ facilitates integrating the system into a more complex model without drawing parallel connections.
@@
Figure 2: Hierarchical specification of an electro-magneto hydrodynamics (EMHD) model [11]: The fluid model f shown in (b) includes the kinetic energy system kin depicted in (a), as well as four additional models not detailed here: an internal energy system int, a thermal conduction model th, and volume and shear viscosity models vol and shr. The electro-magneto hydrodynamics model emhd shown in (c) includes the fluid model f depicted in (b), as well as three additional models not detailed here: an electromagnetic system em, an electric conduction model el, and a model for the electro-kinetic coupling ekc.
Figure 2 shows the hierarchical specification of a plasma model. For simplicity, a spatial domain with isolated boundary is assumed such that only the in-domain interconnection of the different models needs to be dealt with. A red filling indicates that the respective input is a primitive system modeling an irreversible process. A yellow filling signifies that the input is a nested composite system, specified in terms of its own pattern. Recursively substituting into all yellow boxes the respective patterns would yield an equivalent model specification without the hierarchical decomposition. The composition of patterns through substitution is uniquely defined whenever interfaces match (possibly up to a given renaming of the ports). No additional data or handling of ‘edge cases’ is required. One can hence easily add levels of abstraction by refactoring subsystems into reusable parts. In the fluid model, one could for instance factor out the kinetic energy system filling box $\mathtt{kin}$ and the internal energy system filling box $\mathtt{int}$ into a reusable model of an ideal compressible fluid, see [11].
@@ 1.3 Why?
EPHS provide a formal and intuitive modeling language suited for humans and software tools alike. As shown above, increasingly complex models are assembled from simpler and ultimately primitive parts. This fosters the reusability of models and encourages the conceptual separation of different levels of detail. Beyond the modeling process, this can be beneficial when the language is used in interdisciplinary communication and education. Some modeling errors can be completely avoided due to the built-in thermodynamic consistency and the typed interfaces (given by ports with fixed physical quantities). Since interconnection patterns are essentially energy/exergy flow diagrams, EPHS models are particularly useful for thermodynamic analysis and optimization. The compositional nature of EPHS could enable a scalable approach to model transformations, such as discretizations. It could also enable a practical integration of first-principles-based thermodynamic modeling and scientific machine learning.
1.4 Nonequilibrium thermodynamics in a nutshell
@@
Most physical systems are characterized by a huge number of degrees of freedom but keeping track of them is usually neither possible nor useful. Nonequilibrium thermodynamics aims to provide a framework for coarse-grained modeling, meaning that fast dynamics at small scales are not resolved, and a quantity called entropy is introduced to account for the missing information, see, e.g., [6]. Coarse-graining is possible because, despite fluctuations at the microscopic scale, one can eventually observe non-fluctuating properties at a more macroscopic scale; at least this is the case for isolated systems. For instance, temperature is a property that only appears at more macroscopic scales. Thermodynamic properties are classified as extensive when they are proportional to the size of the system (e.g. entropy, volume, mass) or as intensive when they do not depend on the size of the system (e.g. temperature, pressure, chemical potential), see, e.g., [12]
@@
A system is said to be in local thermodynamic equilibrium if for every material point there exists a neighborhood large enough to observe thermodynamic properties, yet small enough for the properties to be essentially non-varying inside the neighborhood, see, e.g., [13]. The size of these neighborhoods corresponds to a macroscopic scale at which the dynamics of the thermodynamic properties can be observed and also modeled.
@@
The first law of thermodynamics restates a fact known from mechanics, namely that energy is a conserved quantity. However, instead of considering directly the mechanical and electromagnetic energy associated to microscopic degrees of freedom, a thermodynamic model with local equilibrium resorts to a phenomenological description. This so-called internal energy is considered to depend on a number of extensive properties, including entropy, and possibly further mechanical or electromagnetic state variables which are observable on the considered macroscopic scale. These quantities then serve as the state variables of a coarse-grained description. Moreover, temperature is defined as the derivative of the internal energy with respect to entropy. Similarly, pressure is given by the derivative with respect to volume and the chemical potential is the derivative with respect to mass, see, e.g., [12].
@@
Models in classical mechanics and electromagnetism exhibit reversible dynamics, meaning that one could in principle arrive again at the initial condition by formally reversing the direction of time (although for many systems this would be practically impossible due to chaos). Incomplete knowledge about the microscopic state immediately leads to an irreversible dynamics and an irreversible dynamics in turn leads to a growth of entropy, as information (e.g. about the initial state of the system) is lost as the system evolves. This is reflected by the second law of thermodynamics which states that the entropy of an isolated system never decreases and that the system evolves towards a state of maximum entropy, see, e.g., [13]. In other words, the uncertainty about the microscopic state can only grow, not decrease.
@@
For models with local equilibrium, the irreversible dynamics can be expressed in terms of so-called thermodynamic fluxes that determine instantaneous changes of coarse-grained state variables. The fluxes are expressed as a function of so-called thermodynamic forces, which are given by local differences in intensive quantities, see, e.g., [13, 14]. For instance, temperature differences cause heat flux, according to Fourier’s law, and differences in chemical potential cause a diffusive mass flux, according to Fick’s law. Further, one may observe cross effects such as a diffusive mass flux caused by temperature differences (Soret effect) and likewise a heat flux caused by differences in chemical potential (Dufour effect). Onsager’s reciprocal relations state that the relation between fluxes and forces must possess a certain symmetry [15]. Specifically, the ratio of the mass flux caused by the Soret effect and the respective temperature difference is equal to the ratio of the heat flux caused by the Dufour effect and the respective difference in chemical potential.
@@ 1.5 Related work
Figure 3: Informal Venn diagram illustrating that EPHS integrates ideas from graphical and energy-based modeling of physical systems with bond graphs (see, e.g., [7, 8]), the metriplectic/GENERIC framework for nonequilibrium thermodynamics (see, e.g., [5, 13]), port-Hamiltonian theory for open, dissipative systems (PHS) (see, e.g., [4]), and applied category theory (ACT) research on the formalization of graphical languages and compositional dynamical systems (see, e.g., [9, 16, 17, 18, 19]).
1.5.1 Bond graphs
Bond graphs provide a graphical notation for representing multi-domain models, particularly those involving mechanical and electrical systems, see, e.g., [7, 8]. The main idea is that elements interact via so-called bonds that represent links for energy exchange. Each bond is associated with two dynamic variables, called flow and effort, whose product gives the exchanged power. Bonds can be connected through two types of junctions: 00-junctions, where the flow variables of all connected ports balance while all effort variables are equal, and $1$-junctions, where the effort variables balance while all flow variables are equal. In both cases, the net power at a junction is zero. The elements are commonly generalized capacitors, inductors, transformers, resistors, etc. The junctions are then seen as generalized Kirchhoff’s circuit laws.
@@ 1.5.2 Port-Hamiltonian systems
Port-Hamiltonian Systems (PHS) were originally inspired by relating the dynamical systems obtained from generalized bond graphs111 While bond graphs commonly use generalized capacitors and inductors as storage elements, so-called generalized bond graphs use only generalized capacitors together with gyrator elements [21]. to Hamiltonian mechanics [20]. While many authors make use of PHS in input-state-output form, PHS generally constitute an acausal systems theory, also inspired by Willems’ behavioral approach [22, 23]. The dynamics of networks composed of energy storing elements and reversible coupling elements (generalized transformers and gyrators) can, from a geometric perspective, be formalized as Hamiltonian systems whose Poisson structure reflects the network topology [24]. The generalization to systems with algebraic constraints and/or external port variables is achieved using Dirac structures [25]. These offer a relational approach that unifies Poisson and presymplectic structures (see, e.g., [1, 26]). By modeling resistive elements as relations imposed on external port variables, Hamiltonian dynamics can be combined with (dissipative) gradient dynamics. A key feature of PHS is their composability: interconnecting simpler systems yields a more complex one. This is achieved by composing the Dirac structures of the individual systems with an ‘interconnecting Dirac structure’ (see, e.g., [4]). Another key feature is that every PHS satisfies a power-balance equation, stating that at every instant the sum of the stored energy and the dissipated energy equals the supplied energy. As the dissipated energy cannot be negative, PHS are inherently passive. In this context, the term ‘energy’ stems from the ‘Hamiltonian’ storage function. For EPHS, this function is interpreted as the stored exergy, making the power-balance equation a statement about rates of exergy transfer. This interpretation fully reconciles the inherent passivity of PHS with the first and second law of thermodynamics, as exergy is indeed dissipated (or destroyed) by irreversible dynamics.
1.5.3 Metriplectic and GENERIC framework
Metriplectic systems as well as the general equation for reversible-irreversible coupling (GENERIC) refer to a framework for combining reversible and irreversible dynamics in a thermodynamically consistent manner, see, e.g. [5, 27, 28, 29, 30]. The reversible part is modeled as a Hamiltonian dynamics, which inherently conserves energy, while the irreversible part is modeled as a (generalized) gradient flow, which inherently drives the system toward maximum entropy. Two additional conditions ensure that the reversible dynamics conserves entropy, and the irreversible dynamics conserves energy. As a result, the combined dynamics respects both the first law and the second law (see, e.g., [13, 6]). While many well-known physical models can be identified as metriplectic systems or instances of the GENERIC, the framework is particularly useful as a guide for developing new thermodynamic models.